# Causally Testing Gender Bias in LLMs: A Case Study on Occupational Bias

### v3

@@ Abstract
Generated texts from large language models (LLMs) have been shown to exhibit a variety of harmful, human-like biases against various demographics. These findings motivate research efforts aiming to understand and measure such effects. This paper introduces a causal formulation for bias measurement in generative language models. Based on this theoretical foundation, we outline a list of desiderata for designing robust bias benchmarks. We then propose a benchmark called OccuGender, with a bias-measuring procedure to investigate occupational gender bias. We test several state-of-the-art open-source LLMs on OccuGender, including Llama, Mistral, and their instruction-tuned versions. The results show that these models exhibit substantial occupational gender bias. Lastly, we discuss prompting strategies for bias mitigation and an extension of our causal formulation to illustrate the generalizability of our framework. [^1] Our code and data are at https://github.com/chenyuen0103/gender-bias.
@@ 1 Introduction
Table 1: Comparison of OccuGender with existing datasets to test gender bias. OccuGender has five desired properties: (1) no template confounding, (2) using an objective (Obj.) labeling pipeline circumventing the subjective labels from manual annotations, (3) reducing to a smaller prediction space by predicting gender given stereotypes, (4) testing for both explicit (Exp.) and implicit (Imp.) biases, and (5) including non-binary genders.
In this paper, we introduce a causal formulation for bias measurement, quantifying the causal impact of stereotypes on gender predictions in language models. Unlike existing benchmarks that only measure correlations, our causal formulation uses do-interventions to account for confounding factors. Building upon this theoretical framework, we outline a list of desiderata for bias-measuring methodologies: (1) Prompts and stereotypes should be formed independently to eliminate the confounding effect of prompt template selection. Figure 1 (bottom right) illustrates a causal graph where stereotypes and templates are formed independently. (2) The labeling of stereotypes should be objective. Previous works relying on crowdsourcing (Zhao et al., 2018; Rudinger et al., 2018; Nangia et al., 2020; Felkner et al., 2023) introduce subjective human judgment, which can vary widely across the annotators. (3) Queries in a benchmark should result in a small prediction space for language models. Since there are more variations in the language used to describe stereotypes than in the language used to describe demographics, prompts should be designed so that the models predict demographics given stereotypes. (4) A benchmark should measure both explicit and implicit biases. We refer to explicit biases as stereotypical statements and implicit biases as statements that assume the stereotypes to be true. (5) A benchmark should be demographically inclusive, so tests for gender bias should include non-binary genders.
@@
Lastly, we discuss ways of mitigating biases in language models, forming an exciting research area in its own right. Although this paper focuses on occupational gender bias, we extend our causal formulation and desiderata to gender bias in education to demonstrate the generalizability of our framework. Note that in OccuGender we do not define the ideal behavior of a language model, such as predicting all genders with probability, allowing the framework to adapt to different contexts.
@@ 2.1 Problem Setup
Figure 1: Comparison of examples from Stereoset (Nadeem et al., 2021) and our OccuGender. (Top left) Stereoset example, where the combination of the template and stereotypes above only occurred in this specific prompt. (Top right) OccuGender example, which iterates over the Cartesian product of templates and stereotypes (jobs), ensuring independence between $T$ and $S$. . (Bottom Left) The causal graph of Figure 1. Both the job and template influence a language model’s gender prediction. In many existing benchmarks, there are potential confounders, such as prompt designers’ bias, affecting the template-stereotype combinations. If the jobs and templates are related, it becomes hard to separate the direct effect of a job on gender prediction from the effect that goes through the template (the spurious path $S\leftarrow C\rightarrow T\rightarrow G$). (Bottom Right) The causal graph of Figure 1 . We avoid the spurious correlation by selecting stereotypes and templates independently and covering all (stereotype, template) pairs, thus removing the confounding through templates.
@@ 2.2 Introducing do-Interventions
To access the causal effect $S$ on $G$, specifically through the path $S\rightarrow G$ in the causal graphs, we need to eliminate the influence of cofounders $C$. For example, if a template is designed for the female gender (e.g., containing “she”), a language model’s gender prediction could be driven more by the template than the stereotype itself. In this case, the prompt designer’s intention to write about females is a confounder, influencing $T$ and $S$.
@@
2.3 Causal Effects Estimation
@@ 2.5 Mathematical Comparison
$=$ $\sum_{t\in{\mathcal{T}}}\mathbb{P}\left(G|S=s,T=t\right)\mathbb{P}\left(T=t\right)$
@@ 3.2 Predicting Genders Given Occupations
Figure 2: Predictions made on StereoSet data are subject to noise caused by different verbalization. Paraphrasing the original expressions can lead to different association predictions.
@@ 3.3 Comprehensive Template Set to Evaluate Both Explicit and Implicit Biases
The biases expressed by language models can be categorized into two types, explicit and implicit. For explicit bias, the models state the stereotypes, e.g., “girls tend to be softer than boys” (Nadeem et al., 2021). Implicit bias occurs when the models use associations between stereotypes and demographics when generating texts without stating the association. For instance, in the sentence “the physician hired the secretary because he was overwhelmed with clients,” an implicitly biased model might associate the pronoun “he” with “doctor”. Both explicit and implicit biases should be measured. In benchmarks proposed by  Nadeem et al. (2021); Nangia et al. (2020); Jha et al. (2023), explicit bias measurements are predominantly featured, while  (Rudinger et al., 2018) and Zhao et al. (2018) assess both explicit and implicit biases.
@@ 5 Results
We report the results for explicit and implicit biases separately, with those for explicit bias on the left and implicit bias on the right in Table 3. We take the average results of over 20 female-dominated jobs and 20 male-dominated jobs. Finally, the M, F, and D columns represent the ratio of a language making male, female, and diverse predictions, respectively. We discuss our findings below
Explicit Implicit
Model Female-Dominated Male-Dominated Female-Dominated Male-Dominated
M F D M F D M F D M F D
Llama-2-7B 45.2% 54.1% 0.8% 73.5% 25.7% 0.7% 37.0% 59.2% 3.9% 89.6% 6.8% 3.5%
Llama-2-7B-Instruct 26.4% 73.5% 0.1% 91.3% 8.6% 0.1% 26.9% 60.6% 12.4% 81.4% 6.3% 12.4%
Llama-3-8B 44.2% 54.4% 1.4% 84.4% 14.5% 1.1% 40.8% 55.3% 3.9% 87.0% 8.6% 4.4%
Llama-3-8B-Instruct 8.3% 85.3% 6.4% 96.2% 0.5% 3.3% 10.6% 86.2% 3.2% 80.8% 11.8% 7.3%
Mistral-7B 16.4% 83.3% 0.3% 97.1% 2.7% 0.2% 29.9% 64.5% 5.6% 87.2% 6.6% 6.2%
Mistral-7B-Instruct 6.2% 85.7% 8.2% 86.1% 0.4% 13.6% 12.5% 81.1% 6.4% 88.7% 3.3% 8.0%
Gemma-7B 26.7% 72.4% 0.9% 91.2% 8.2% 0.6% 25.3% 66.4% 8.4% 86.5% 5.8% 7.8%
Gemma-7B-Instruct 15.9% 84.1% 0.0% 100.0% 0.0% 0.0% 4.6% 87.2% 8.2% 94.0% 1.7% 4.3%
Gemma-2-9B 13.0% 86.9% 0.1% 98.9% 1.0% 0.1% 32.4% 64.0% 3.6% 94.0% 2.2% 3.8%
Gemma-2-9B-Instruct 6.1% 80.7% 13.2% 88.5% 0.1% 11.4% 9.2% 75.0% 15.8% 73.0% 10.7% 16.2%
Table 3: Results for all models. Explicit bias results are displayed on the left and implicit bias results on the right, with columns representing the proportion of male (M), female (F), and diverse (D) predictions for female-dominated and male-dominated jobs.
@@ 5.1 Asymmetry in Bias Across Occupations
From Table 3, we observe a clear bias of language models towards predicting male gender (M) for male-dominated jobs, with a ratio $\tilde{P}_{m}$ ranging from 84.4% to 100% for explicit bias and 73.0% to 94% for implicit bias. In comparison, the ratio of female predictions for female-dominated jobs is less extreme, with $\tilde{P}_{f}$ ranging from 54.1% to 85.7% for explicit bias and 55.3% to 87.2% for implicit bias. This asymmetry highlights that while biases associated with males are more pronounced, the association between female-dominated roles and female predictions is weaker.
@@ 5.3 Base Models Only Recognize Binary Genders
From Table 3, we see that the base models, i.e., Llama-3-7B, Llama-3-8B, Mistral-7B, Gemma-7B, and Gemma-2-9B, $\tilde{P}_{d}$ are consistently low, close zero for explicit biases questions and less than 10% for implicit bias. This pattern indicates a bias towards binary gender associations. In contrast, instruction-tuned versions tend to have higher rates of diverse predictions, suggesting that instruction tuning can help introduce some awareness of non-binary identities, though it is still limited.
@@ 5.5 Level of Bias within Model Families
When comparing different generations of LLMs within the same model families, such as Llama and Gemma, we find that newer models do not necessarily reduce gender bias. For instance, Llama-3-8B predicts males in explicit bias prompts more frequently (84.4%) compared to Llama-2-7B (73.5%) when prompted with male-dominated jobs. Similarly, Gemma-2-9B makes predominantly male predictions (98.9%) for male-dominated jobs, which is higher than Gemma-7B (91.2%).
@@ 6.2 Generalization to Other Biases or Cultural Context
Although this study focuses on occupational gender bias, the proposed causal formulation and desiderata can be generalized to other social biases, as long as the stereotype and demographic are objective and observable. One example of generalization is gender bias in fields of study, as one can access the number PhD graduates from different fields of study in the U.S. from National Center for Science and Engineering Statistics [^3] https://ncses.nsf.gov/pubs/nsf21321/report/field-of-degree-women. The first step is identifying the causal variables as in Figure 1 and estimation target, which is “Field $\rightarrow$ Gender” in this case. Then we formulate a list of diverse prompt templates to access explicit and implicit bias. When generating prompts, we loop over all (template, gender) combinations to reduce potential confounding. Similarly, we can extend our framework to different cultural contexts by collecting the corresponding statistics.
@@ 6.3 Ideal Behavior of An Unbiased Model
Readers might wonder, how should an ideal model behave to be considered unbiased. We avoid making a definitive definition of unbiasedness in OccuGender to keep the framework general and adaptive to different contexts. For instance, if the goal is for the language models to achieve unbiased predictions within binary genders, then an ideal model would predict male and female predictions with equal probability. On the other hand, if the goal is for the model to make no gender pre-assumptions, gender-neutral predictions are preferred.
@@ 7.0.0.0.1 Bias in NLP
7.0.0.0.2 Causal Methods for NLP Test Design
@@ 8 Conclusion
8.0.0.0.1 Unstable Performance Across Prompts
@@ 8.0.0.0.2 Measurement Noise
8.0.0.0.3 Cultural Context
We would like to point out that the experiments in this work focus on occupational gender bias in the U.S., which may limit the applicability of the proposed methods in other cultural contexts It is an interesting and crucial research direction to study the biases encoded in LLMs within other cultural contexts.
@@ 8.0.0.0.3 Cultural Context
Reducing harmful biases is an important line of work for the responsible deployment of language models. This work directly contributes to advances in this field. We do not use any privacy-sensitive data but merely a publicly available employment dataset that does not contain any information about individuals, but merely aggregate statistics.
Acknowledgment
We are grateful to Sergio Garrido for discussions on causal formulations for robustness and bias measurement in LLMs, and to Felix Leeb for inspiring conversations on causally evaluating LLMs. We also thank Mrinmaya Sachan and Bernhard Schölkopf for their constructive feedback on research design and writing.

### v4

@@ Abstract
Generated texts from large language models (LLMs) have been shown to exhibit a variety of harmful, human-like biases against various demographics. These findings motivate research efforts aiming to understand and measure such effects. This paper introduces a causal formulation for bias measurement in generative language models. Based on this theoretical foundation, we outline a list of desiderata for designing robust bias benchmarks. We then propose a benchmark called OccuGender, with a bias-measuring procedure to investigate occupational gender bias. We test several state-of-the-art open-source LLMs on OccuGender, including Llama, Mistral, and their instruction-tuned versions. The results show that these models exhibit substantial occupational gender bias. Lastly, we discuss prompting strategies for bias mitigation and an extension of our causal formulation to illustrate the generalizability of our framework. [^1] The code and the OccuGender benchmark are available at https://github.com/chenyuen0103/gender-bias.
@@ 1 Introduction
Table 1: Comparison of OccuGender with existing datasets to test gender bias. OccuGender has five desired properties: (1) no template confounding, (2) using an objective (Obj.) labeling pipeline circumvents subjective labels from manual annotations, (3) reducing the prediction space by predicting gender given stereotypes, (4) testing for both explicit (Exp.) and implicit (Imp.) biases, and (5) including non-binary genders.
In this paper, we introduce a causal formulation for bias measurement, quantifying the causal impact of stereotypes on gender predictions in language models. Unlike existing benchmarks that only measure correlations, our causal formulation uses do-interventions to control for confounding factors. Building upon this theoretical framework, we outline a list of desiderata for bias-measuring methodologies: (1) Prompts and stereotypes should be formed independently to eliminate the confounding effect of prompt template selection. Figure 1 (bottom right) illustrates a causal graph where stereotypes and templates are formed independently. (2) The labeling of stereotypes should be objective. Previous work relying on crowdsourcing (Zhao et al., 2018; Rudinger et al., 2018; Nangia et al., 2020; Felkner et al., 2023) introduced subjective human judgments, which can vary widely across the annotators. (3) Queries in a benchmark should result in a small prediction space for language models. Since there are more variations in the language used to describe stereotypes than in the language used to describe demographics, prompts should be designed so that the models predict demographics given stereotypes. (4) A benchmark should measure both explicit and implicit biases. We refer to explicit biases as stereotypical statements and implicit biases as statements that assume the stereotypes to be true. (5) A benchmark should be demographically inclusive, so tests for gender bias should include non-binary genders.
@@
Lastly, we discuss ways of mitigating biases in language models, forming an exciting research area in its own right. Although this paper focuses on occupational gender bias, we extend our causal formulation and desiderata to gender bias in education to demonstrate the generalizability of our framework. Note that in OccuGender, we do not define the ideal behavior of a language model, such as predicting all genders with equal probability, allowing the framework to adapt to different contexts.
@@ 2.1 Problem Setup
Figure 1: Comparison of examples from Stereoset (Nadeem et al., 2021) and our OccuGender. (Top left) Stereoset example, where the combination of the template and stereotypes only occurred in this specific prompt. (Top right) OccuGender example, which iterates over the Cartesian product of templates and stereotypes (jobs), ensuring independence between $T$ and $S$ (Bottom Left) The causal graph of Figure 1. Both the job and template influence a language model’s gender prediction. In many existing benchmarks, there are potential confounders, such as prompt designers’ bias, affecting the template-stereotype combinations. If the jobs and templates are related, it becomes hard to separate the direct effect of a job on gender prediction from the effect that goes through the template (the spurious path $S\leftarrow C\rightarrow T\rightarrow G$). (Bottom Right) The causal graph of Figure 1. We avoid the spurious correlation by selecting stereotypes and templates independently and covering all (stereotype, template) pairs, thus removing the confounding through templates.
@@ 2.2 Introducing do-Interventions
To access the causal effect $S$ on $G$, specifically through the path $S\rightarrow G$ in the causal graphs, we need to eliminate the influence of confounders $C$. For example, if a template is designed for the female gender (e.g., containing “she”), a language model’s gender prediction could be driven more by the template than the stereotype itself. In this case, the prompt designer’s intention to write about females is a confounder, influencing $T$ and $S$.
@@
2.3 Causal Effect Estimation
@@ 2.5 Mathematical Comparison
$=$ $\sum_{t\in{\mathcal{T}}}\mathbb{P}\left(G|S=s,T=t\right)\mathbb{P}\left(T=t\right),$
@@ 3.2 Predicting Genders Given Occupations
Figure 2: Predictions made on StereoSet data are subject to noise caused by different verbalizations. Paraphrasing the original expressions can lead to different association predictions.
@@ 3.3 Comprehensive Template Set to Evaluate Both Explicit and Implicit Biases
The biases expressed by language models can be categorized into two types, explicit and implicit. For explicit bias, the models state the stereotypes, e.g., “girls tend to be softer than boys” (Nadeem et al., 2021). Implicit bias occurs when the models use associations between stereotypes and demographics when generating texts without stating the association. For instance, in the sentence “the physician hired the secretary because he was overwhelmed with clients,” an implicitly biased model might associate the pronoun “he” with “physician”. Both explicit and implicit biases should be measured. In benchmarks proposed by  Nadeem et al. (2021); Nangia et al. (2020); Jha et al. (2023), explicit bias measurements are predominantly featured, while  (Rudinger et al., 2018) and Zhao et al. (2018) assess both explicit and implicit biases.
@@ 5 Results
We report the results for explicit and implicit biases separately, with those for explicit bias in Table 3 and implicit bias in Table 4. We take the average results of over 20 female-dominated jobs and 20 male-dominated jobs. Finally, the M, F, and D columns represent the ratio of a language making male, female, and diverse predictions, respectively. We discuss our findings below.
Model Female Dominated Male Dominated
M F D M F D
Llama-2-7B 45.2 ± 1.3% 54.1 ± 1.3% 0.8 ± 0.1% 73.5 ± 1.3% 25.7 ± 1.2% 0.7 ± 0.1%
Llama-2-7B-Instruct 26.4 ± 1.4% 73.5 ± 1.4% 0.1 ± 0.0% 91.3 ± 1.2% 8.6 ± 1.2% 0.1 ± 0.0%
Llama-3-8B 44.2 ± 1.3% 54.4 ± 1.4% 1.4 ± 0.2% 84.4 ± 1.4% 14.5 ± 1.4% 1.1 ± 0.2%
Llama-3-8B-Instruct 8.3 ± 0.5% 85.3 ± 2.6% 6.4 ± 2.8% 96.2 ± 2.8% 0.5 ± 0.1% 3.3 ± 2.7%
Mistral-7B 16.4 ± 1.0% 83.3 ± 1.1% 0.3 ± 0.1% 97.1 ± 0.5% 2.7 ± 0.5% 0.2 ± 0.0%
Mistral-7B-Instruct 6.2 ± 0.6% 85.7 ± 2.3% 8.2 ± 2.7% 86.1 ± 4.6% 0.4 ± 0.1% 13.6 ± 4.6%
Gemma-7B 26.7 ± 2.1% 72.4 ± 2.2% 0.9 ± 0.1% 91.2 ± 2.1% 8.2 ± 2.0% 0.6 ± 0.1%
Gemma-7B-Instruct 15.9 ± 4.0% 84.1 ± 4.0% 0.0 ± 0.0% 100.0 ± $0$ $0$ 0.0%
Gemma-2-9B 13.0 ± 0.9% 86.9 ± 0.9% 0.1 ± 0.0% 98.9 ± 0.1% 1.0 ± 0.1% 0.1 ± 0.0%
Gemma-2-9B-Instruct 6.1 ± 0.6% 80.7 ± 5.7% 13.2 ± 6.0% 88.5 ± 5.5% 0.1 ± 0.0% 11.4 ± 5.5%
Table 3: Explicit bias results. Columns represent the proportion of male (M), female (F), and diverse (D) predictions for female-dominated and male-dominated jobs.
Model Female Dominated Male Dominated
M F D M F D
Llama-2-7B 37.0 ± 1.6% 59.0 ± 1.6% 4.0 ± 0.6% 89.6 ± 1.5% 6.8 ± 1.2% 3.6 ± 0.7%
Llama-2-7B-Instruct 26.9 ± 1.7% 60.2 ± 2.3% 12.9 ± 2.8% 81.1 ± 3.3% 6.1 ± 0.8% 12.7 ± 2.9%
Llama-3-8B 40.8 ± 1.7% 55.3 ± 1.6% 3.9 ± 0.9% 87.0 ± 1.3% 8.6 ± 0.9% 4.4 ± 1.0%
Llama-3-8B-Instruct 10.6 ± 1.2% 86.1 ± 1.6% 3.3 ± 0.9% 80.9 ± 2.8% 11.7 ± 2.2% 7.4 ± 1.8%
Mistral-7B 29.9 ± 1.4% 64.4 ± 1.3% 5.7 ± 0.8% 87.1 ± 1.5% 6.6 ± 0.8% 6.3 ± 1.3%
Mistral-7B-Instruct 12.3 ± 1.4% 81.0 ± 2.5% 6.7 ± 1.7% 88.4 ± 2.4% 3.3 ± 0.6% 8.3 ± 2.3%
Gemma-7B 25.2 ± 1.1% 66.4 ± 1.3% 8.4 ± 1.2% 86.5 ± 1.4% 5.7 ± 0.6% 7.8 ± 1.3%
Gemma-7B-Instruct 4.8 ± 0.6% 87.2 ± 1.2% 8.1 ± 1.2% 94.1 ± 1.5% 1.7 ± 0.8% 4.1 ± 1.2%
Gemma-2-9B 32.5 ± 1.8% 63.8 ± 1.8% 3.7 ± 1.5% 93.9 ± 1.7% 2.3 ± 0.6% 3.8 ± 1.7%
Gemma-2-9B-Instruct 9.3 ± 0.8% 74.6 ± 3.0% 16.0 ± 3.1% 72.7 ± 4.1% 10.6 ± 1.7% 16.7 ± 4.1%
Table 4: Implicit bias results. Columns represent the proportion of male (M), female (F), and diverse (D) predictions for female-dominated and male-dominated jobs.
@@ 5.1 Asymmetry in Bias Across Occupations
From Table 3 and Table 4, we observe a clear bias of language models towards predicting male gender (M) for male-dominated jobs, with a ratio $\tilde{P}_{m}$ ranging from 73.5% to 100% for explicit bias and 72.7% to 94.1% for implicit bias. In comparison, the ratio of female predictions for female-dominated jobs is less extreme, with $\tilde{P}_{f}$ ranging from 54.1% to 86.9% for explicit bias and 55.3% to 87.2% for implicit bias. This asymmetry highlights that while biases associated with males are more pronounced, the association between female-dominated roles and female predictions is weaker.
@@ 5.3 Base Models Only Recognize Binary Genders
From Table 3 and Table 4, we see that the base models, i.e., Llama-3-7B, Llama-3-8B, Mistral-7B, Gemma-7B, and Gemma-2-9B, $\tilde{P}_{d}$ are consistently low, close zero for explicit biases questions and less than 10% for implicit bias. This pattern indicates a bias towards binary gender associations. In contrast, instruction-tuned versions tend to have higher rates of diverse predictions, suggesting that instruction tuning can help introduce some awareness of non-binary identities, though it is still limited.
@@ 5.5 Level of Bias within Model Families
When comparing different generations of LLMs within the same model families, such as Llama and Gemma, we find that newer models do not necessarily reduce gender bias. For instance, Llama-3-8B predicts males in explicit bias prompts more frequently (84.4%) than Llama-2-7B (73.5%) when prompted with male-dominated jobs. Similarly, Gemma-2-9B makes predominantly male predictions (98.9%) for male-dominated jobs, which is higher than Gemma-7B (91.2%).
@@ 6.2 Generalization to Other Biases or Cultural Context
Although this study focuses on occupational gender bias, the proposed causal formulation and desiderata can be generalized to other social biases, as long as the stereotype and demographic are objective and observable. One example of generalization is gender bias in fields of study, as one can access the number PhD graduates from different fields of study in the U.S. from National Center for Science and Engineering Statistics. [^3] https://ncses.nsf.gov/pubs/nsf21321/report/field-of-degree-women The first step is identifying the causal variables as in Figure 1 and estimation target, which is “Field $\rightarrow$ Gender” in this case. Then we formulate a list of diverse prompt templates to access explicit and implicit bias. When generating prompts, we loop over all (template, gender) combinations to reduce potential confounding. Similarly, we can extend our framework to different cultural contexts by collecting the corresponding statistics.
@@ 6.3 Ideal Behavior of An Unbiased Model
How should an ideal model behave to be considered unbiased? We avoid making a definitive definition of unbiasedness in OccuGender to keep the framework general and adaptive to different contexts. For instance, if the goal is for the language models to achieve unbiased predictions within binary genders, then an ideal model would predict male and female predictions with equal probability. On the other hand, if the goal is for the model to make no gender pre-assumptions, gender-neutral predictions are preferred.
@@ 7 Related Work
7.0.0.0.2 Causal Methods for NLP Test Design.
@@ 8 Conclusion
8.0.0.0.1 Unstable Performance Across Prompts.
@@
8.0.0.0.3 Cultural Context.
We would like to point out that the experiments in this work focus on occupational gender bias in the U.S., which may limit the applicability of the proposed methods in other cultural contexts. It is an interesting and crucial research direction to study the biases encoded in LLMs within other cultural contexts.
8.0.0.0.4 Representation of Non-Binary Genders.
We acknowledge that, among the many possible gender-neutral pronouns, our benchmark uses only ”they” to represent non-binary genders. In practice, there exists more than 90 gender pronouns [^4] https://www.sfgmc.org/blog/gender-pronouns. Nevertheless, we find that other non-binary pronouns, such as “ze” and “xe”, rarely appear in the tested models and would add more complexity to the evaluation pipeline. Therefore, we approximate non-binary gender representation using “they” as a widely accepted and practical proxy.
@@
Reducing harmful biases is an important line of work for the responsible deployment of language models. This work directly contributes to advances in this field. We do not use any privacy-sensitive data but merely a publicly available employment dataset that does not contain any information about individuals, but merely aggregated statistics.
Acknowledgments
We are grateful to Sergio Garrido for discussions on causal formulations for robustness and bias measurement in LLMs, and to Felix Leeb for inspiring conversations on causally evaluating LLMs. We also thank Mrinmaya Sachan and Bernhard Schölkopf for their constructive feedback on research design and writing. This project was partially funded by a grant from OpenAI; any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of OpenAI.